# GMIMS-DRAGONS: A Faraday Depth Survey of the Northern Sky Covering 350 to 1030 MHz

### v1

@@ 7.2 Comparison with GMIMS-LBS
The strong agreement between GMIMS-LBS and DRAGONS allows for the maps to be combined into a single, full-sky image across the overlapping frequency range. We show examples of such a combined map in Figure 37, where we have averaged the maps in the overlapping declination range. This highlights the potential of the full GMIMS initiative in allowing for polarization and Faraday depth mapping of large-scale patterns simultaneously covering both northern and southern Galactic hemispheres. Examples of forthcoming studies that take advantage of this are: (i) a follow-up to the analysis of the asymmetry between the northern and southern Galactic latitudes explored in Dickey et al. (2022) using GMIMS-HBN and STAPS (C. van Bergen et al., 2026, in preparation) and (ii) an analysis of the large-scale magnetic field reversal using STAPS and the higher frequencies (>500>500 MHz) of DRAGONS (R. Booth et al., 2025, in preparation). Other future work will be able to benefit from the high Faraday depth resolution provided by GMIMS-LBS and the lower frequencies of DRAGONS to investigate Faraday complexity in the more nearby Galactic volume by using only the lower end of the DRAGONS band (<500<500 MHz).
7.3 Perseus-Taurus Bubble
Tahani et al. (2022a) further reconstructed the 3D magnetic field vector of the Perseus molecular cloud, revealing that another structure beyond the Per-Tau bubble must have interacted with the Perseus cloud and shaped its field lines. They called this feature Per2, a finding subsequently confirmed by the kinematic studies of (Kounkel et al., 2022). Faraday depth maps can be used not only to study the magnetic field morphology of the Per-Tau bubble and its surrounding environment, but also to help identify the Per2 structure, its shape, and the extent of its impact (M. Tahani et al., 2026, in preparation). The bubble stands out as a depolarized region in the DRAGONS M0 map (top panel of Figure 38) and its outline, along with features within the bubble, appears across multiple Faraday depth values (examples shown the bottom two panels of Figure 38).
Figure 38: The Per-Tau bubble in DRAGONS M0 (top), PI at Faraday depth -5.0 rad m-2 (middle), and PI at 5.0 rad m-2 (bottom). The white dashed circle is the boundary of the bubble given by Bialy et al. (2021). The masked (grey) region is Tau A, with the same masking as in Figure 3235.
8 Conclusions and Future Work
We have used the DRAO-15 telescope to produce the DRAGONS polarization survey, which comprises the single-antenna, low-frequency, northern hemisphere component of GMIMS. This survey included a science commissioning phase for the DRAO-15, since this was the first science initiative with this instrument. The DRAGONS survey is spatially Nyquist sampled, covering a declination range of $20\arcdeg$ to $90\arcdeg$, and a frequency range of 350–1030 MHz. The 83 kHz sampling allowed for effective RFI excision prior to binning into 1 MHz channelization for the final data cubes. We used the GPS-based ALBUS software to correct for ionospheric Faraday rotation effects in the Stokes $Q$ and $U$ data, which can be significant at the lower frequencies of DRAGONS, particularly near solar maximum. The observing strategy of $360\arcdeg$ azimuth scans over the course of seven months allowed for adequate baseline correction and ground subtraction, while the residual differences between the resulting duplicated maps of the sky (East and West) provided a position-dependent estimate of the errors in the maps.
@@ 8 Conclusions and Future Work
The DRAGONS cubes provide an immense improvement over the Dwingeloo data (Brouw & Spoelstra, 1976), the best previously available polarization data in this frequency range, in terms of both spatial and spectral sampling and coverage. This allows for detailed studies of Faraday complexity and large-scale patterns in Faraday rotation across the entire northern sky. We observe Faraday complexity in the form of multiple Faraday depth peaks and broadened structures in over half of the observed region. While the Faraday depth first moment, M1, map reveals many similarities with the M1 maps of other GMIMS component surveys, the structures along the Faraday depth axis will undoubtedly provide critical information for probing the LOS component of the Galactic magnetic field.
@@
There are already several projects in progress making use of the DRAGONS dataset, including modeling of the large-scale Galactic magnetic field reversal (R. Booth et al., 2025, in preparation), a comparison of Faraday depth spectra and moments maps between DRAGONS and LoTSS DR2, and efforts to produce $20\arcmin$ angular resolution polarization maps by combining DRAGONS with CHIME data in 400–800 MHz. In combination with other GMIMS datasets, DRAGONS provides an exciting opportunity to investigate the effects of depolarization and the polarization horizon by studying Faraday depth spectra produced using different frequency sub-bands. DRAGONS will also be useful for total intensity, Stokes $I$ studies. In the frequency range covered by DRAGONS, the PyGDSM total intensity model is constrained by only the Haslam et al. (1982) 408 MHz data and the Dwingeloo 820 MHz data (Berkhuijsen, 1971). The near-complete frequency sampling of DRAGONS may allow for more detailed spectral index mapping in total intensity. The DRAGONS polarization data cube in conjunction with this could also be used to separate out the synchrotron component.
@@ Acknowledgments
The telescope is a prototype developed through an international collaboration for the Square Kilometre Array (SKA) project. We acknowledge the crucial contributions of the US SKA Technology Development Project (TDP), supported by the National Science Foundation (NSF), for the antenna’s advanced optical design, high-speed mechanical drive system, and stable tower structure. The instrument’s excellent performance also relies on the Band 1 feed, developed by Chalmers University of Technology at the Onsala Space Observatory (OSO) as part of their contribution to the SKA.
@@
We also thank Bryan Gaensler, Gary Hinshaw, Lewis Knee, Roland Kothes, Josh MacEachern, Wasim Raja, Michael Rupen, Ken Tapping, Alec Thomson, and Cameron Van Eck for their constructive feedback over the course of the project, and we thank Paeton Hughes for helping us with data quality assessment.
@@
A.O. was partly supported by the Dunlap Institute at the University of Toronto. R.A.B. was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) Vanier scholarship and the University of Calgary Izaak Walton Killam Doctoral Scholarship. This research has been supported by NSERC Discovery Grants (PIs: J.C.B., T.L.L., and A.S.H). G.M. was supported by an NSERC Undergraduate Student Research Award, and L.M.C. was supported by an International Undergraduate Research Award funded by UBC Okanagan’s International Student Initiative and delivered in partnership with the Irving K. Barber Faculty of Science.
@@
FITS files containing the data summarized in Table 4 are available through the Canadian Advanced Network for Astronomical Research (CANFAR), doi: to be published.

### v2

@@ VII.3 Perseus-Taurus Bubble
Tahani et al. (2022b) further reconstructed the 3D magnetic field vector of the Perseus molecular cloud, revealing that another structure beyond the Per-Tau bubble must have interacted with the Perseus cloud and shaped its field lines. They called this feature Per2, a finding subsequently confirmed by the kinematic studies of (Kounkel et al., 2022). Faraday depth maps can be used not only to study the magnetic field morphology of the Per-Tau bubble and its surrounding environment, but also to help identify the Per2 structure, its shape, and the extent of its impact (M. Tahani et al., 2026, in preparation). The bubble stands out as a depolarized region in the DRAGONS M0 map (top panel of Figure 38) and its outline, along with features within the bubble, appears across multiple Faraday depth values (examples shown the bottom two panels of Figure 38).
Figure 38: The Per-Tau bubble in DRAGONS M0 (top), PI at Faraday depth -5.0 rad m-2 (middle), and PI at 5.0 rad m-2 (bottom). The white dashed circle is the boundary of the bubble given by Bialy et al. (2021). The masked (gray) region is Tau A, with the same masking as in Figure 3235.
VIII Conclusions and Future Work
We have used the DRAO-15 telescope to produce the DRAGONS polarization survey, which comprises the single-antenna, low-frequency, northern-hemisphere component of GMIMS. This survey included a science commissioning phase for the DRAO-15, since this was the first science initiative with this instrument. The DRAGONS survey is spatially Nyquist sampled, covering a declination range of $20\arcdeg$ to $90\arcdeg$, and a frequency range of 350–1030 MHz. The 83 kHz sampling allowed for effective RFI excision prior to binning into 1 MHz channelization for the final data cubes. We used the GPS-based ALBUS software to correct for ionospheric Faraday rotation effects in the Stokes $Q$ and $U$ data, which can be significant at the lower frequencies of DRAGONS, particularly near solar maximum. The observing strategy of $360\arcdeg$ azimuth scans over the course of seven months allowed for adequate baseline correction and ground subtraction, while the residual differences between the resulting duplicated maps of the sky (east and west) provided a position-dependent estimate of the errors in the maps.
@@ VIII Conclusions and Future Work
The DRAGONS cubes provide an immense improvement over the Dwingeloo data (Brouw and Spoelstra, 1976), the best previously available polarization data in this frequency range, in terms of both spatial and spectral sampling and coverage. This allows for detailed studies of Faraday complexity and large-scale patterns in Faraday rotation across the entire northern sky. We observe Faraday complexity in the form of multiple Faraday depth peaks and broadened structures in over half of the observed region. While the Faraday depth first moment, M1, map reveals many similarities with the M1 maps of other GMIMS component surveys, the structures along the Faraday depth axis will undoubtedly provide critical information for probing the LOS component of the Galactic magnetic field.
@@
There are already several projects completed or in progress that make use of the DRAGONS dataset, including modeling of the large-scale Galactic magnetic field reversal (Booth et al., 2026), a comparison of Faraday depth spectra and moments maps between DRAGONS and LoTSS DR2, and efforts to produce $20\arcmin$ angular resolution polarization maps by combining DRAGONS with CHIME data in 400–800 MHz. In combination with other GMIMS datasets, DRAGONS provides an exciting opportunity to investigate the effects of depolarization and the polarization horizon by studying Faraday depth spectra produced using different frequency sub-bands. DRAGONS will also be useful for total intensity, Stokes $I$ studies. In the frequency range covered by DRAGONS, the PyGDSM total intensity model is constrained by only the Haslam et al. (1982) 408 MHz data and the Dwingeloo 820 MHz data (Berkhuijsen, 1971). The near-complete frequency sampling of DRAGONS may allow for more detailed spectral index mapping in total intensity. The DRAGONS polarization data cube in conjunction with this could also be used to separate out the synchrotron component.
@@ Acknowledgments
The telescope is a prototype developed through an international collaboration for the SKA project. We acknowledge the crucial contributions of the US SKA Technology Development Project (TDP), supported by the National Science Foundation (NSF), for the antenna’s advanced optical design, high-speed mechanical drive system, and stable tower structure. The instrument’s excellent performance also relies on the Band 1 feed, developed by Chalmers University of Technology at the Onsala Space Observatory (OSO) as part of their contribution to the SKA.
@@
We also thank Bryan Gaensler, Gary Hinshaw, Lewis Knee, Roland Kothes, Josh MacEachern, Wasim Raja, Michael Rupen, Ken Tapping, Alec Thomson, and Cameron Van Eck for their constructive feedback over the course of the project, and we thank Paeton Hughes for helping us with data quality assessment. We appreciate the comments from the anonymous referee, which helped clarify some details of the data processing.
@@
A.O. was partly supported by the Dunlap Institute at the University of Toronto. The Dunlap Institute is funded through an endowment established by the David Dunlap family and the University of Toronto. R.A.B. was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) Vanier scholarship and the University of Calgary Izaak Walton Killam Doctoral Scholarship. This research has been supported by NSERC Discovery Grants (PIs: J.C.B., T.L.L., and A.S.H). G.M. was supported by an NSERC Undergraduate Student Research Award, and L.M.C. was supported by an International Undergraduate Research Award funded by UBC Okanagan’s International Student Initiative and delivered in partnership with the Irving K. Barber Faculty of Science.
@@
FITS files containing the data summarized in Table 4 are available through the Canadian Advanced Network for Astronomical Research (CANFAR), doi: 10.11570/25.0104.